In a beauty contest, the ages of the contestants are normally distributed with a mean of


24 years and a standard deviation of 1.5. What percent of the contestants have age


between 22.5 and 27 years old?

1
Expert's answer
2022-04-06T17:15:34-0400

"P(22.5<X<27)=P(\frac{22.5-24}{1.5}<Z<\frac{27-24}{1.5})=P(-1<Z<2)=0.9772-0.1587=0.8185"